Read the paragraph and answer the questions:
The sliding frictional force between a 4kg box and the floor is 15N. It is pushed across the floor with a constant force such that it accelerates at 0.8m/s^2.
Ques 1:What is the force applied to the box?
Ques 2:If this same force calculated is applied to the box which is now placed on a frictionless surface,what is the new acceleration produced?

Grade:8

1 Answers

Ravleen Kaur
6 years ago
1.
Force Appilied = Friction force + mass x acceleration
= 15 + 4 X 0.8
= 18.2 N
2.
Accleration = Force / mass = 18.2/4 = 4.55 m/s2

regards,

IHE